John's age is four times the age of Jake. If the difference between their age is 63, how old is John and how old is Jake?
Answer by ankor@dixie-net.com(22740)

let x = John's present age
let y = Jake's
:
Write an equation for each statement
:
John's age is four times the age of Jake.
x = 4y
If the difference between their age is 63,
x - y = 63
how old is John and how old is Jake?
In the above equation, replace x with 4y (from the 1st statement)
4y - y = 63
3y = 63
y = 63/3
y = 21 is Jake's age
then
x = 4(21)
x = 84 yrs is John's age
;
:
Check: 84 - 21 = 63